# Portable Biomechanics Laboratory: Clinically Accessible Movement Analysis from a Handheld Smartphone

### v1

@@ Abstract
Background The way a person moves is a direct reflection of their neurological and musculoskeletal health, yet it remains one of the most underutilized vital signs in clinical practice. Although clinicians visually observe movement impairments, they lack accessible and validated methods to objectively measure movement in routine care. This gap prevents wider use of biomechanical measurements in practice, which could enable more sensitive outcome measures or earlier identification of impairment.
Methods In this work, we present our Portable Biomechanics Laboratory (PBL), which includes a secure, cloud-enabled smartphone app for data collection and a novel algorithm for fitting biomechanical models to this data. We extensively validated PBL’s biomechanical measures using a large, clinically representative and heterogeneous dataset with synchronous ground truth. Next, we tested the usability and utility of our system in both a neurosurgery and sports medicine clinic.
Results We found joint angle errors within 3 degrees and pelvis translation errors within several centimeters across participants with neurological injury, lower-limb prosthesis users, pediatric inpatients, and controls. In addition to being easy and quick to use, gait metrics computed from the PBL showed high reliability (ICCs > 0.9) and were sensitive to clinical differences. For example, in individuals undergoing decompression surgery for cervical myelopathy, the modified Japanese Orthopedic Association (mJOA) score is a common patient-reported outcome measure; we found that PBL gait metrics not only correlated with mJOA scores but also demonstrated greater responsiveness to surgical intervention than the patient-reported outcomes.
Conclusions These findings support the use of handheld smartphone video as a scalable, low-burden tool for capturing clinically meaningful biomechanical data, offering a promising path toward remote, accessible monitoring of mobility impairments in clinical populations. To promote further research and clinical translation, we release the first method for measuring whole-body kinematics from handheld smartphone video validated in clinical populations: https://IntelligentSensingAndRehabilitation.github.io/MonocularBiomechanics/.
keywords:
Computer Vision | AI Gait Analysis | Biomechanical Modeling
@@ Introduction
Many clinical conditions have pronounced movement phenotypes which are rarely measured in clinical practice (1, 2, 3). For example, temporal parameters, including stride time and swing time, have been shown to predict fall risk in older adults (4). Patients with knee osteoarthritis (KOA) also show altered gait kinematics, which can predict disease progression (5, 6). Gait kinematics can also characterize recovery dynamics after neurological injury and even predict responses to particular interventions (7, 8). The Stroke Recovery and Rehabilitation Roundtable emphasized the importance of incorporating movement quantification into clinical trials, while noting that logistical challenges in capturing such data remain a significant barrier (9). Traditionally, gait analysis requires a specialized laboratory equipped with an optical motion capture (OMC) system and force plates, making it expensive, time-consuming, and typically only covered by insurance in limited circumstances (10). In some cases, such as orthopedic surgical planning, force plate measurements and electromyography are required to ensure the precision of marker-based measurements.
@@
Currently, gait is more commonly characterized using the 10-meter walk or 6-minute walk tests, which capture only speed and endurance. Other assessments—such as the Timed Up and Go (TUG) (11), the Berg Balance Scale (BBS) (12), and the Functional Gait Assessment (FGA) (13)—evaluate movement quality using clinical scoring methods that are either subjective or limited to stopwatch-based timing, but do not directly quantify the movement itself. Scalable, accessible movement analysis in the clinic would substantially improve upon the status quo.
@@
Recent advances in AI-based methods offer a promising path forward. Multi-view markerless motion capture (MMMC) requires much less time to acquire and less manual post-processing than OMC. In recent years, numerous studies have demonstrated that MMMC produces similar kinematics to OMC in multiple populations (14, 15, 16, 17, 18, 19, 20, 21, 22, 23). OpenCap (24) even allows MMMC to be performed using videos collected from multiple smartphones mounted on tripods. We have also developed methods for MMMC (25, 26, 27, 25, 28). A key differentiator of our approaches is the use of MuJoCo (29), a high-performance physics simulator for machine learning that supports differentiable biomechanical models. This enables us to directly end-to-end optimize the kinematics and skeleton scale from videos. Our end-to-end approach outperforms alternatives that reproduce two-stage marker-based pipelines that first estimate virtual marker locations and then compute inverse kinematics on these trajectories (26, 27). While MMMC is faster than OMC, having multiple synchronized cameras that are spatially calibrated is much less clinically accessible than simply recording with a smartphone.
To make movement analysis more accessible, we previously developed a smartphone app for collecting smartphone videos from clinic (30), which we call our Portable Biomechanics Laboratory (PBL), and validated that we could accurately detect gait events from this data (31). We also tested whether our end-to-end approach to biomechanics can reconstruct kinematics from these smartphone videos and wearable sensors (32). In addition to demonstrating that we could fuse this data, we also found that when the participant is not obscured, we can obtain accurate knee kinematics from smartphone video alone. Fig. 1 provides an overview of our approach. In this study, we validate the accuracy of whole-body kinematics computed with our system and test it in the clinic, showing that the resulting gait metrics are sensitive to the clinical history of our participants. Next, we compare the responsiveness of gait metrics to traditional outcome measures. Specifically, we monitored the gait of a cohort of participants undergoing decompression surgery for cervical myelopathy (CM). We find that gait metrics show higher responsiveness than the modified Japanese Orthopedic Association (mJOA), a commonly used outcome measure for CM (33, 34). Our study demonstrates that quantitative movement analysis can be integrated into a clinic in an accessible and usable workflow, producing accurate and useful results, and that it can be more responsive than traditional clinical outcomes assessments.
Figure 1: Methods Overview. We introduce a method for biomechanically grounded movement analysis in clinical settings using a handheld smartphone. A) Researchers held a smartphone (optionally with gimbal) while following a participant walking. Our system has no specific requirements regarding viewing angle, distance to subject, or therapist assistance. B) Recorded smartphone video and optional wearable sensor data are stored in the cloud, and processed using PosePipe, an open-source package implementing computer vision models for person tracking and keypoint detection. C) To reconstruct movement, we represent movement as a function that outputs joint angles, which—combined with body scaling parameters and evaluated through forward kinematics—generate a posed biomechanical model in 3D space. This untrained model is compared to video-extracted joint locations and optionally smartphone sensor data to compute a loss. This loss guides backpropagation to iteratively refine both the kinematic trajectory and body scale. D) Initially, the representation lacks knowledge of the person’s movements and scale (e.g., height, limb proportions), but after optimization, it typically tracks joint locations within 15 mm in 3D and 5 pixels in 2D.
@@ Methods
We developed a platform called Portable Biomechanics Laboratory (PBL), which records color video, optional depth video, and internal phone sensor information (accelerometer, gyroscopes, and orientation) from a moving, handheld smartphone (Fig. 1A). The application can optionally be connected to wearable IMUs with all modalities closely synchronized (30, 32). After a recording is finished, PBL uploads the data to a secure cloud platform where data can be pulled into PosePipe (35), an open-source package we developed for managing and processing videos.
@@
Using PosePipe (35), we identified the subject of interest in each video using DeepSortYOLOv4 (36). Next, we extracted 2D and 3D virtual marker locations, termed "keypoints", from each video frame using MeTRAbs-ACAE (37) (Fig. 1B). We previously identified the 87 keypoints from the MoVi dataset (38) to be sufficient for full-body kinematics (39).
@@
We obtained kinematics by optimizing the fit of biomechanical models using our end-to-end differentiable biomechanics framework (32, 26). A complete description is in the Supplemental Methods; here we provide a brief overview (Fig. 1C,D). We represented a movement trajectory as a neural network that maps time to pelvis location and joint angles, termed an "implicit function". This network was optimized to minimize error between predicted virtual markers and those extracted from video, using GPU-accelerated physics simulation in MuJoCo (29, 40).
@@
We applied our monocular fitting approach to three datasets. The first of these recorded clinical participants with the PBL system and also an MMMC system. The second used iPhone videos and an OMC system, but only contained able-bodied individuals. The third dataset reflects our intended use case of videos recorded from participants seen in various clinics, which we used to demonstrate our method can capture clinically meaningful features of gait. Combined, these datasets contain 21 hours of data.
@@
    Our MMMC dataset includes 11.7 hours of recordings from 161 sessions with 148 participants performing walking and lower-limb functional tasks. Of these, 121 participants were recorded simultaneously using both the PBL system (30) and our MMMC system (41, 26). This subset included 33 healthy controls, 6 pediatric inpatients, 48 lower limb prosthesis users (LLPUs), and 40 participants with neurological injury. An additional 27 neurological inpatients were recorded using only the MMMC system on a clinical floor. For neurologic patients, we extracted clinical outcome scores—such as the 10 Meter Walk Test (10MWT) and Berg Balance Scale (BBS) (12)—from 54 participants across 95 visits.
@@
    The OMC dataset used was BML-MoVi (38), a publicly available dataset of 90 healthy controls performing a variety of sports and upper/lower body movements. It contains 3.8 hours of synchronized video (captured with an iPhone 7) and OMC data using a dense 87-marker set. We fit this dataset using our end-to-end differentiable biomechanics approach (26), applying the same biomechanical model used for the MMMC and PBL datasets.
@@
    Our in-clinic dataset includes 5.5 hours of video from 74 participants recorded with the PBL system during outpatient clinic visits: 55 from a neurosurgery clinic for cervical myelopathy (CM) and 19 from a clinic for knee osteoarthritis (KOA). After their standard physician consultations, participants walked in the clinic hallway while a researcher recorded them using the PBL system. CM participants also completed the Modified Japanese Orthopedic Association (mJOA) questionnaire (33, 34), which scores upper and lower extremity function. Seventeen CM participants were recorded at visits both before and after surgery.
@@
To evaluate the accuracy of the smartphone-based fits compared to OMC and MMMC, we computed the Median Joint Angle Error (MJAE). This is the median of the absolute difference between the two systems over a trial for each joint.
@@
To compare moving (handheld) vs static camera video, we also computed fits from single cameras in the MMMC dataset. As the MMMC videos captured participants from every angle, we examined the effect of camera viewpoints on joint angle error by classifying the camera position relative to the person into front, back, ipsilateral and contralateral.
@@
For static camera and some PBL fits, we report the Root Translation Error (RTE), defined as the average Euclidean error between the pelvis location from the PBL reconstruction and ground truth OMC/MMMC reconstruction. For PBL, we computed the RTE only for fits with camera rotation but no translation (e.g. during a TUG trial).
When aggregating errors over trials, we report the median and normalized interquartile range (nIQR) for MJAE where $nIQR=IQR\cdot 0.7413$.
Gait Metrics and Sensitivity to Clinical Condition
@@
To test whether our portable system collected clinically meaningful information, we calculated the Gait Deviation Index (GDI) (42). Briefly, the GDI uses a dimensionality reduction technique (42, 43) on cycle-aligned joint kinematics and then measures the distance in this subspace to a distribution of normative gait trials. For the normative distribution, we used PBL and MMMC reconstruction from controls in the MMMC dataset as well as an additional 136 participants without gait impairments recruited at the American Society of Biomechanics meeting, resulting in 9,008 normative steps. We also computed double support time and cadence using a model for gait event detection we previously developed and validated (31, 30).
Clinical metrics must be reliable, valid, and responsive. To assess reliability, we computed intra-class correlations (ICCs) between steps taken by the same individual on the same day (44). Construct validity of the GDI was tested by comparing groups with known differences (e.g., transfemoral vs. transtibial amputees) and by correlating kinematic metrics with established clinical scores, including the mJOA, 10MWT, and BBS. To assess responsiveness, we compared the smartphone-based metrics and the mJOA following surgical decompression in the CM dataset using the Standardized Response Mean (SRM) (45, 46). Confidence intervals on SRMs were computed using a jackknife resampling procedure (47, 46).

### v2

@@ Abstract
Movement directly reflects neurological and musculoskeletal health, yet objective biomechanical assessment is rarely available in routine care. We introduce Portable Biomechanics Laboratory (PBL), a secure platform for fitting biomechanical models to video collected with a handheld, moving, smartphone. We validate this approach on over 15 hours of data synchronized to ground truth motion capture, finding mean joint‑angle errors < 3° and pelvis‑translation errors of a few centimeters across patients with neurological‑injury, lower‑limb prosthesis users, pediatric in‑patients, and controls. In > 5 hours of prospective deployments to neurosurgery and sports‑medicine clinics, PBL was easy to setup, yielded highly reliable gait metrics (ICC > 0.9), and detected clinically relevant differences. For cervical‑myelopathy patients, its measurement of gait quality correlated with modified Japanese Orthopedic Association (mJOA) scores and were responsive to clinical intervention. Handheld smartphone video can therefore deliver accurate, scalable, and low‑burden biomechanical measurement, enabling greatly increased monitoring of movement impairments. We release the first clinically-validated method for measuring whole-body kinematics from handheld smartphone video at https://IntelligentSensingAndRehabilitation.github.io/MonocularBiomechanics/.
@@ Introduction
Many clinical conditions have pronounced movement phenotypes which are rarely measured in clinical practice [jacquelin2010gait, balaban_gait_2014, morris_biomechanics_2001]. For example, temporal parameters, including stride time and swing time, have been shown to predict fall risk in older adults [hamacher_kinematic_2011]. Patients with knee osteoarthritis (KOA) also show altered gait kinematics, which can predict disease progression [heiden_knee_2009, sosdian_longitudinal_2014]. Gait kinematics can also characterize recovery dynamics after neurological injury and even predict responses to particular interventions [chow_longitudinal_2021, awad_these_2020]. The Stroke Recovery and Rehabilitation Roundtable emphasized the importance of incorporating movement quantification into clinical trials, while noting that logistical challenges in capturing such data remain a significant barrier [kwakkel_standardized_2017]. Traditionally, gait analysis requires a specialized laboratory equipped with an optical motion capture (OMC) system and force plates, making it expensive, time-consuming, and typically only covered by insurance in limited circumstances [abdullah_multibody_2024]. In some cases, such as orthopedic surgical planning, force plate measurements and electromyography are also required to ensure the precision of marker-based measurements.
@@
Currently, gait is more commonly characterized using the 10-meter walk or 6-minute walk tests, which capture only speed and endurance. Other assessments—such as the Timed Up and Go (TUG) [podsiadlo_timed_1991], the Berg Balance Scale (BBS) [berg_clinical_1992], and the Functional Gait Assessment (FGA) [wrisley_reliability_2004]—evaluate movement quality using clinical scoring methods that are either subjective or limited to stopwatch-based timing, but do not directly quantify the movement itself. Scalable, accessible movement analysis in the clinic would substantially improve upon the status quo.
@@
Recent advances in AI-based methods offer a promising path forward. Multi-view markerless motion capture (MMMC) requires much less time to acquire and less manual post-processing than OMC. In recent years, numerous studies have demonstrated that MMMC produces similar kinematics to OMC in multiple populations [kanko_concurrent_2021, kanko_inter-session_2021, kanko_assessment_2021, riazati_absolute_2022, song_markerless_2023, wren_comparison_2023, horsak_repeatability_2024, dsouza_comparison_2024, wang_evaluation_2025, min_biomechanical_2024]. We have developed methods for MMMC [cotton_biomechanical_2025, cotton_differentiable_2025, firouzabadi_biomechanical_2024, unger_differentiable_2025] using MuJoCo [todorov_mujoco_2012], a high-performance physics simulator for machine learning that supports differentiable biomechanical models. This enables us to directly end-to-end optimize the kinematics and skeleton scale from videos. Our end-to-end approach outperforms alternatives that reproduce multi-stage marker-based pipelines that first estimate virtual marker locations and then compute inverse kinematics on these trajectories [cotton_differentiable_2025, firouzabadi_biomechanical_2024]. Although MMMC is faster than OMC, installing and calibrating several high-performance cameras remains a significant barrier to clinical implementation.
OpenCap substantially lowers the technical burden of MMMC by using two or three tripod-mounted smartphones, achieving < 5 degrees of joint-angle error relative to OMC [uhlrich_opencap_2023] and detecting movement patterns that classify disease [ruth_video-based_2025, min_biomechanical_2024]. While OpenCap is a significant and widely used contribution, the requirement for two tripods placed at ±45 degrees is not always feasible in clinical settings, which often include small exam rooms and tight hallways. In these contexts, a single handheld camera offers a more practical alternative. Monocular video analysis has been used clinically with people with Parkinson’s disease [acevedo_visionmd_2025], pediatric patients with cerebral palsy [segado_data-driven_2025], and stroke [lonini_video-based_2022]. However, these approaches infer 2D or 3D joint locations, or non-biomechanically-grounded mesh representations, rather than biomechanical joint angles used in traditional movement analysis. Recent methods [koleini_biopose_2025, xia_reconstructing_2025] have explored data-driven biomechanical reconstruction from a single camera, but have not been trained or validated on clinical populations. To enable data capture in real-world clinical settings, we previously developed a smartphone application for collecting handheld videos in clinics [cimorelli_validation_2024], which we refer to as the Portable Biomechanics Laboratory (PBL), and validated gait event detection from these recordings [cotton_transforming_2022]. We subsequently demonstrated our end-to-end biomechanical pipeline can reconstruct biomechanically-grounded kinematics from handheld smartphone video, with or without wearable sensors, achieving accurate knee kinematics from video alone [peiffer_fusing_2024].
Figure 1: Study Design Overview. We first validated the Portable Biomechanics Laboratory (PBL) against an optical motion capture dataset in which control participants performed a range of upper- and lower-limb activities. We then evaluated PBL in clinical populations by recording participants performing lower-limb tasks using both multi-camera markerless motion capture and PBL in hospital laboratories and on clinical floors. Finally, we deployed PBL in clinic hallways and therapy gyms, recording participants after clinical visits to assess the relationship between PBL-derived gait quality scores and standard clinical measures.
The central goal of the present study is to evaluate whether this smartphone-based biomechanics pipeline can produce accurate kinematics across diverse patient populations, movement tasks, and clinical environments that show construct validity against clinical metrics. We perform this validation by first evaluating our system against OMC in control populations, next against MMMC in clinical populations, and finally, our intended use case, collect data in actual clinic hallways to compare against clinical scores (Fig. 1). Overall, we find PBL has high joint-angle accuracy compared to OMC and MMMC, concordance with clinical scores, and is responsive to clinical intervention. Together, these evaluations demonstrate that quantitative movement analysis can be integrated into routine clinical workflows using a single handheld smartphone for accurate and clinically relevant biomechanical measurement.
@@ Methods
We developed a platform called Portable Biomechanics Laboratory (PBL), which records color video, optional depth video, and internal phone sensor information (accelerometer, gyroscopes, and orientation) from a moving, handheld smartphone (Fig. 2A). The application can optionally be connected to wearable IMUs with all modalities closely synchronized [cimorelli_validation_2024, peiffer_fusing_2024]. After a recording is finished, PBL uploads the data to a secure cloud platform where data can be pulled into PosePipe [cotton_posepipe_2022], an open-source package we developed for managing and processing videos.
Figure 2: Biomechanical Reconstruction Overview. We introduce a method for biomechanically grounded movement analysis in clinical settings using a handheld smartphone. A) Researchers held a smartphone (optionally with gimbal) while following a participant walking. Our system has no specific requirements regarding viewing angle, distance to subject, or therapist assistance. B) Recorded smartphone video and optional wearable sensor data are stored in the cloud, and processed using PosePipe, an open-source package implementing computer vision models for person tracking and keypoint detection. C) To reconstruct movement, we represent movement as a function that outputs joint angles, which—combined with body scaling parameters and evaluated through forward kinematics—generate a posed biomechanical model in 3D space. This untrained model is compared to video-extracted joint locations and optionally smartphone sensor data to compute a loss. This loss guides backpropagation to iteratively refine both the kinematic trajectory and body scale. D) Initially, the representation lacks knowledge of the person’s movements and scale (e.g., height, limb proportions), but after optimization, it typically tracks joint locations within 15 mm in 3D and 5 pixels in 2D.
@@
Using PosePipe [cotton_posepipe_2022], we identified the subject of interest in each video using DeepSortYOLOv4 [wojke_simple_2017]. Next, we extracted 2D and 3D virtual marker locations, termed "keypoints", from each video frame using MeTRAbs-ACAE [sarandi_learning_2023] (Fig. 2B). We previously identified the 87 keypoints from the MoVi dataset [ghorbani_movi_2021] to be sufficient for full-body kinematics [cotton_optimizing_2023].
@@
We obtained kinematics by optimizing the fit of biomechanical models using our end-to-end differentiable biomechanics framework [peiffer_fusing_2024, cotton_differentiable_2025]. We note that this study estimated joint kinematics only; no inverse dynamics or kinetic estimation was performed. A complete description is in the Supplemental Methods; here we provide a brief overview (Fig. 2C,D). We represented a movement trajectory as a neural network that maps time to pelvis location and joint angles, termed an "implicit function". This network was optimized to minimize error between predicted virtual markers and those extracted from video using GPU-accelerated physics simulation in MuJoCo [todorov_mujoco_2012, caggiano_myosuite_2022].
Run Time
Preprocessing a 6 second video takes 10 seconds on a server-grade GPU, 1 minute on a laptop-grade GPU and scales linearly with video length. Biomechanical fitting takes 1 minute on a server-grade GPU, <5 minutes on a laptop-grade GPU, and does not scale with video length.
@@
We applied our monocular fitting approach to three datasets (Fig. 1). The first is a publicly available dataset that used iPhone videos and an OMC system, but only contained able-bodied individuals. The second of these recorded clinical participants with the PBL system and also an MMMC system. The third dataset reflects our intended use case of videos recorded from participants seen in various clinics, which we used to demonstrate our method can capture clinically meaningful features of gait. Combined, these datasets contain 21 hours of data.
@@
    The OMC dataset used was BML-MoVi [ghorbani_movi_2021], a publicly available dataset of 90 healthy controls performing a variety of sports and upper/lower body movements. It contains 3.8 hours of synchronized video (captured with an iPhone 7) and OMC data using a dense 87-marker set. We fit both modalities of this dataset using our end-to-end differentiable biomechanics approach [cotton_differentiable_2025], applying the same biomechanical model used for the MMMC and PBL datasets.
@@
    Our MMMC dataset includes 11.7 hours of recordings from 161 sessions with 148 participants performing walking and lower-limb functional tasks. Of these, 121 participants were recorded simultaneously using both the PBL system [cimorelli_validation_2024] and our MMMC system [cotton_markerless_2023, cotton_differentiable_2025]. This subset included 33 healthy controls, 6 pediatric inpatients, 48 lower limb prosthesis users (LLPUs), and 40 participants with neurological injury. An additional 27 neurological inpatients were recorded using only the MMMC system on a clinical floor. For neurologic patients, we extracted clinical outcome scores—such as the 10 Meter Walk Test (10MWT) and Berg Balance Scale (BBS) [berg_clinical_1992]—from 54 participants across 95 visits. For the BBS, we classify a score of $\leq$ 50 as a fall risk [lusardi_determining_2017] and $\leq$ 40 as high fall risk [shumway-cook_predicting_1997].
@@
    Our in-clinic dataset includes 5.5 hours of video from 74 participants recorded with the PBL system during outpatient clinic visits: 55 from a neurosurgery clinic for cervical myelopathy (CM) and 19 from a clinic for knee osteoarthritis (KOA). After their standard physician consultations, participants walked in the clinic hallway while a researcher recorded them using the PBL system. CM participants also completed the Modified Japanese Orthopedic Association (mJOA) questionnaire [kopjar_validity_2011, tetreault_modified_2017], which scores upper and lower extremity function. Seventeen CM participants were recorded at visits both before and after surgery.
@@
To evaluate the accuracy of the smartphone-based fits compared to OMC and MMMC, we computed the Median Absolute Error, defined as the median of joint-angle errors within each trial followed by the median of these per-trial medians across trials. We used Median Absolute Error because both the within-trial and across-trial error distributions were non-Gaussian. When aggregating errors across trials, we report the median and the normalized interquartile range (nIQR), where $nIQR=IQR\cdot 0.7413$.
@@
To compare moving (handheld) vs. static camera video, we also computed monocular fits from single cameras in the MMMC dataset. Because MMMC records participants from all directions, we examined the influence of camera viewpoint on joint-angle error by classifying viewpoint as front, back, ipsilateral, or contralateral.
@@
For static-camera and relevant PBL fits, we report the Root Translation Error (RTE), defined as the average Euclidean error between the pelvis location from the PBL reconstruction and the ground-truth OMC/MMMC reconstruction. For PBL, RTE was computed only for trials involving camera rotation but no camera translation (e.g., TUG trials).
For comparison to OpenCap and other methods, we also compute the Mean Absolute Error following the procedure in [uhlrich_opencap_2023], detailed in the Supplementary Methods.
Clinical Construct Validity of Smartphone-Derived Gait Metrics
Gait Deviation Index
@@
To reduce kinematic timeseries to an interpretable score representing gait quality, we calculated the Gait Deviation Index (GDI) [schwartz_gait_2008]. Briefly, the GDI uses a dimensionality reduction technique [schwartz_gait_2008, marks_measuring_2018] on cycle-aligned joint kinematics and then measures the distance in this subspace to a distribution of normative gait trials (for which we used 9,008 steps of 206 control participants from both monocular and MMMC systems). We computed GDI, measured between 0 and 100 with 100 being normative gait, on every gait cycle for a participant and then averaged GDI over a session. Finally, we also computed double support time (DST) and cadence using a model for gait event detection that we previously developed and validated [cotton_transforming_2022, cimorelli_validation_2024].
Construct Validity
Clinical metrics must be repeatable, valid, and responsive. We assessed the repeatability of kinematics from our system by computing intra-class correlations (ICCs) of the GDI between steps taken by the same individual on the same day [bartko_intraclass_1966]. We tested construct validity of our system using t-tests to separate GDI distributions between groups. To further test construct validity, we also compared the GDI to established clinical scores, including the mJOA, 10MWT, and BBS. To evaluate the responsiveness of smartphone-based gait metrics to surgical intervention, we analyzed longitudinal pre- and post-operative changes in mJOA, cadence, DST, and GDI in a cohort of cervical myelopathy participants (n = 17) that were recorded before and after surgery. Paired changes were first assessed for statistical significance using the Wilcoxon signed-rank test. The magnitude and direction of change were then quantified using the corresponding Wilcoxon effect size, which provides a standardized measure on a [1,1][-1,1] scale [fritz_effect_2012]. Ninety-five percent confidence intervals for the effect sizes were estimated via bootstrapping. In addition, changes in raw knee kinematics before and after surgery were assessed using paired SPM1D t-tests [pataky_generalized_2010, pataky_one-dimensional_2012].